OpenStudy (anonymous):

x-intercept=3, y-intercept=2/5. Find the slope-intercept form?

OpenStudy (anonymous):

\[y=-(15/2)x+c\]

OpenStudy (anonymous):

this looks different from my answer:y=2/5x+3?

OpenStudy (anonymous):

slope intercept form is : \(\large y=mx+b \) since y-intercept is given, then b=2/5 also the line passes through the (3, 0), the x-intercept so \(\large 0=m\cdot 3+\frac{2}{5} \) \(\large 0=m\cdot 3+\frac{2}{5} \rightarrow m=-\frac{2}{15} \) so the equation you want is: \(\large y=-\frac{2}{15}x+\frac{2}{5} \)

OpenStudy (anonymous):

I worked the problem, I get lost with equasions, thank you

OpenStudy (anonymous):

dpalnc's Answer Is Right.Sorry I Replied IN Hurry.

OpenStudy (anonymous):

A mountain has an elevation of 19,609 feet. In 1911, the glacier on the peak covered 5 acres. By 2003, the glacier melted to 1 acre. (A) what is the yearly rate in change acres/year I cannot figure out an equasion for this story problem?